# SpecCLIP: Aligning and Translating Spectroscopic Measurements for Stars

### v3

@@ 6 Summary
  • We explore the use of SBI as an alternative to MLPs for downstream parameter estimation. SBI provides improved uncertainty modeling and higher precision for certain parameters, albeit at greater inference cost and lower model capacity.
@@
Looking ahead, we plan to extend this framework to additional spectroscopic modalities, including LAMOST medium-resolution spectra (MRS, Li et al., 2024a), APOGEE infrared spectra (Majewski et al., 2017), the Subaru PFS spectra (Takada et al., 2014), and DESI DR1 spectra (DESI Collaboration et al., 2025). Our approach can be readily adapted to new instruments by pre-training modality-specific encoders and aligning them with contrastive objectives and decoder structures developed in this work. In future iterations, we will also explore efficient adaptation via neural network adapters, enabling scalable multi-survey alignment with minimal computational cost. A large-scale application and catalog release are planned in forthcoming work.
@@ Acknowledgements
The code and relevant data required to reproduce the results in this paper will be made publicly available upon acceptance of the manuscript. Detailed instructions and links will be provided in the final version.
@@
Table 5: Similar to Table 1 for “LRS Models”, but the results for all CLIP-based models are from the alignment between the LAMOST LRS MT and Gaia XP MT (see Section 5.6 for discussion, and Appendices D and E for model details).
@@
Figure 7: Similar to Figure 2, but the results for all CLIP-based models are from the alignment between the LAMOST LRS MT and Gaia XP MT (see Section 5.6 for discussion, and Appendices D and E for model details).
Appendix A Additional Results for Elemental-abundance Estimation
Figure 6 presents an example of chemical-abundance predictions from one input spectrum. This result is obtained using SBI trained with a single model for all parameters simultaneously. The inset in the upper right corner shows results from simulation-based calibration (SBC). Most elemental abundances exhibit well-calibrated posteriors, except for [Ti/Fe] and [Ca/Fe], which fall below the 0.05 significance threshold. This provides an example where, despite the model learns posterior distributions well overall, further tuning of the SBI architecture or its hyperparameters is necessary to achieve reliable multivariate inference of elemental abundances–an aspect we did not further refine in this work.
In Table 5, we provide downstream task results using CLIP-based models aligned between the LAMOST LRS MT and Gaia XP MT encoders. This serves as a comparison to Table 1, where the alignment is performed between the LAMOST LRS MT and Gaia XP OAE models. We find that the latter configuration shows overall comparable or better performance across the evaluation metrics, possibly due to the stronger pretraining of the XP OAE model. For completeness, we include in Figure 7 an external comparison analogous to Figure 2 in the main text, but using the models from Table 5.
@@ Appendix B Continuum Fitting Algorithm
Although we apply continuum fitting only to the blue segment ($4000\,\text{\AA }\leq\lambda\leq 5600\,\text{\AA }$) of the LAMOST LRS spectra for our analysis, we describe here the full continuum-fitting algorithm, which is designed to robustly estimate the stellar continuum over the full wavelength range ($3850\,\text{\AA }\leq\lambda\leq 9000\,\text{\AA }$), and remains effective under varying SNRs. The method takes as input the observed wavelength array $\mathbf{w}$, the corresponding flux array $\mathbf{f}$, and an estimate of the average SNR, and returns a smooth continuum model $\mathbf{c}$.
@@
    Red Segment Correction: A fourth-order polynomial fit is iteratively applied to the red-side segment. At each step, outliers deviating more than $3\sigma$ below the fit (or any point below the fit) are replaced by the polynomial value. This process is repeated up to 8 iterations to suppress absorption features and stabilize the continuum estimate.
@@ Appendix C Normalizing Flows for Parameter Estimation
Normalizing flows provide a flexible and tractable approach to model complex probability distributions by transforming a simple base distribution through a sequence of invertible and differentiable mappings. In SBI, we use normalizing flows to learn an approximate posterior $q_{\phi}(\bm{\theta}\mid\mathbf{x})$, conditioned on observations $\mathbf{x}$, following the Neural Posterior Estimation (NPE) framework.
@@
$q_{\phi}(\bm{\theta}\mid\mathbf{x})=p_{Z}(f^{-1}_{\phi}(\bm{\theta};\mathbf{x}))\left|\det\left(\frac{\partial f^{-1}_{\phi}}{\partial\bm{\theta}}\right)\right|=p_{Z}(\mathbf{z})\left|\det\left(\frac{\partial f_{\phi}}{\partial\mathbf{z}}\right)\right|^{-1},$ (C1)
@@ C.1 Masked Autoregressive Flow (MAF)
MAF (Papamakarios et al., 2017)) models the forward transformation $\mathbf{z}\mapsto\bm{\theta}$ as a sequence of autoregressive operations. Each component of $\bm{\theta}$ is computed as:
$\theta_{i}=\mu_{i}(\bm{\theta}_{<i};\mathbf{x})+\sigma_{i}(\bm{\theta}_{<i};\mathbf{x})\cdot z_{i},$ (C4)
where $\mu_{i}$ and $\sigma_{i}$ are outputs of neural networks that depend on the previous components $\bm{\theta}_{<i}$ and the conditioning variable $\mathbf{x}$. The Jacobian of this transformation is lower triangular, which allows the log-determinant to be computed efficiently:
$\log\left|\det\left(\frac{\partial\bm{\theta}}{\partial\mathbf{z}}\right)\right|=\sum_{i}\log\sigma_{i}(\bm{\theta}_{<i};\mathbf{x}).$ (C5)
@@ C.2 Neural Spline Flow (NSF)
NSF (Durkan et al., 2019) generalizes MAF by replacing affine transformations with monotonic rational-quadratic splines. Each component is transformed as:
$\theta_{i}=\text{spline}_{i}(z_{i};\psi_{i}(\bm{\theta}_{<i},\mathbf{x})),$ (C6)
where $\psi_{i}$ are spline parameters (bin widths, heights, and derivatives) predicted by a neural network conditioned on $\bm{\theta}_{<i}$ and $\mathbf{x}$. The log-determinant of the Jacobian is given by:
@@
where each derivative term is efficiently computed from the spline’s analytical form.
@@ Appendix D Pre-trained Models
In this paper, we tried two different kinds of pre-trained models, one is the transformer-based model and the other is a MLP-based auto-encoder. Both kinds of networks have the same number of trainable parameters (42.7 million). We use a batch size of 64 per GPU. The models are optimized using AdamW with a learning rate of $1\times 10^{-5}$ and a weight decay of 0.1. The learning rate follows a cosine annealing schedule with linear warm-up.
@@ D.1 Transformer-Based Spectral Pre-trained Models
We describe two transformer models designed for masked reconstruction of stellar spectra, based on Parker et al. (2024), each tailored to the characteristics of a different input dataset: LAMOST and Gaia XP.
@@ D.1.1 Masked Spectral Modeling for LAMOST Spectra
The input is linearly embedded and added to learned positional embeddings. Then it is passed through $N=6$ layers of standard transformer blocks with $H=6$ attention heads. The output is normalized and decoded via a linear projection.
@@
To train the model in a self-supervised fashion, we apply a chunk-based masking strategy. Given an input sequence of length $T^{\prime}$, we divide it conceptually into $M=6$ equal-length segments and randomly select a contiguous chunk of width w=10w=10777In previous versions, this number was misstated; the correct value used in the experiments is 10. within each segment to mask. For each chunk, its starting index is sampled uniformly from the allowable range within the segment. This ensures that the masked regions are distributed across the sequence and sufficiently separated.
Formally, let the $i$-th segment span indices $s_{i},s_{i}+\ell]$ where $\ell=\lfloor T^{\prime}/M\rfloor$. Then the masked region for segment $i$ is:
@@
where $m_{t}=1$ if $t$ is masked, and 0 otherwise.
LAMOST spectra benefit from local patterns and detailed features. Thus, using dense chunk-based masking and slicing helps the transformer leverage locality while preserving the global context. We train this model with 8 GPUs over a total of 128 epochs, a process that requires roughly 20 hours.
@@ D.1.2 Masked Spectral Modeling for Gaia XP Spectra
As in the LAMOST model, we standardize the spectrum and pre-pend two scalar tokens:
$x^{\prime}_{0}=\log_{10}\mu,\quad x^{\prime}_{1}=\log_{10}\sigma$ (D5)
@@
We train this model with 8 GPUs and a total of 191 epochs, requiring roughly 40 hours.
@@ D.2 MLP-based Autoencoder
For comparison with the transformer-based models, we also construct an MLP-based autoencoder to pre-train the stellar spectra. For both LAMOST LRS and Gaia XP, the network architectures are similar: an initial projection layer of shape $\text{input\_dim},\text{hidden\_dim}]$; the encoder consists of three MLP blocks, each with structure $\text{hidden\_dim},3\times\text{hidden\_dim},\text{hidden\_dim}]$; followed by a bottleneck layer of shape $\text{hidden\_dim},768]$. The decoder mirrors this structure to reconstruct the input spectra. For LAMOST LRS, we use $\text{input\_dim}=1462+1$ and $\text{hidden\_dim}=1245$; for Gaia XP, we use $\text{input\_dim}=343+2$ and $\text{hidden\_dim}=1290$. The former includes one additional standard deviation as the first token, while the latter includes both the mean and standard deviation as the first two tokens. We train the LRS model on 4 GPUs for 128 epochs, taking approximately 140 minutes, and the Gaia XP model 4 GPUs for 191 epochs, taking approximately 150 minutes.
@@ E.1 Projection Networks
For the projection networks, the LAMOST LRS model follows Parker et al. (2024). We first obtain the output from a pre-trained LAMOST LRS model, then go through a cross-attention module with a learnable query vector. The cross-attention module has four attention heads with an output dimension of 768. Finally, we have a MLP layer with hidden features having $4\times 768$ dimension. We did not compress the information more in this projected network, in order to investigate the information gain without compression. Therefore, the dimension of the final projected embedding is still 768. The number of trainable parameters is about 7.1 million.
@@
Each pathway includes a residual MLP block at the end with hidden dimension $4\times 768$.
@@ E.2 Decoders
For the cross-modal decoder of the CLIP-split model, we takes only the shared representation as input. The decoder architecture depends on the output dimension:
@@ Appendix F Loss Curves
This section presents the loss curve comparisons between the CLIP-pr and CLIP-split models, as shown in Figure 8. While the CLIP-pr model achieves a lower CLIP loss during training, it exhibits lower absolute cosine similarity on test pairs, as shown in Table 2. This discrepancy may arise from two factors. First, the geometry of the high-dimensional embedding spaces—768 dimensions in the CLIP-pr projected embedding versus 512 in the CLIP-split shared embedding—tends to yield lower cosine similarity in the former. Second, the CLIP loss focuses on relative alignment rather than absolute similarity. Additionally, the CLIP-pr model yields slightly lower cross-modal prediction losses. In contrast, the CLIP-split model converges more quickly in learning the reconstruction and achieves marginally lower reconstruction loss, likely due to its design of a non-shared projected embedding space.

### v4

@@ VII Summary
  • We explore the use of SBI as an alternative to MLPs for downstream parameter estimation. SBI provides improved uncertainty modeling and higher precision for certain parameters, albeit at a higher inference cost and a lower model capacity.
@@
Looking ahead, we plan to extend this framework to additional spectroscopic modalities, including LAMOST medium-resolution spectra (MRS, 2024ApJS..273...18L), APOGEE infrared spectra (2017AJ....154...94M), Subaru PFS spectra (2014PASJ...66R...1T), and DESI DR1 spectra (2025arXiv250314745D). Our approach can be readily adapted to new instruments by pre-training modality-specific encoders and aligning them with contrastive objectives and decoder structures developed in this work. In future iterations, we will also explore efficient adaptation via neural network adapters, enabling scalable multi-survey alignment with minimal computational cost. In forthcoming work, a large-scale application and catalog release is planned.
@@ Acknowledgements
All observational data used in this work are publicly available from the archives. A frozen version of the SpecCLIP software used in this analysis has been archived on Zenodo (zhao_2025_17824840) in compliance with the AAS Journals software policy. No proprietary data were used.
[…]
@@
Appendix A Additional Results for Elemental-abundance Estimation
@@
Table 5: Similar to Table 1 for “LRS Models”, but the results for all CLIP-based models are from the alignment between the LAMOST LRS MT and Gaia XP MT (see Section VI.6 for discussion, and Appendices D and E for model details).
@@
Figure 7: Similar to Figure 2, but the results for all CLIP-based models are from the alignment between the LAMOST LRS MT and Gaia XP MT (see Section VI.6 for discussion, and Appendices D and E for model details).
Figure 6 presents an example of chemical-abundance predictions from one input spectrum. This result is obtained using SBI trained with a single model for all parameters simultaneously. The inset in the upper right corner shows results from simulation-based calibration (SBC). Most elemental abundances exhibit well-calibrated posteriors, except for [Ti/Fe] and [Ca/Fe], which fall below the significance threshold of 0.05. This provides an example where, despite the model learns posterior distributions well overall, further tuning of the SBI architecture or its hyper-parameters is necessary to achieve reliable multivariate inference of elemental abundances–an aspect we did not further refine in this work.
In Table 5, we provide downstream task results using CLIP-based models aligned between the LAMOST LRS MT and Gaia XP MT encoders. This serves as a comparison to Table 1, where alignment is performed between the LAMOST LRS MT and Gaia XP OAE models. We find that the latter configuration shows overall comparable or better performance across the evaluation metrics, possibly due to the stronger pretraining of the XP OAE model. For completeness, in Figure 7 we include an external comparison analogous to Figure 2 in the main text, but using the models from Table 5.
@@ Appendix B Continuum Fitting Algorithm
Although we applied continuum fitting only to the blue segment ($4000\,\text{\AA }\leq\lambda\leq 5600\,\text{\AA }$) of the LAMOST LRS spectra for our analysis, we describe here the full continuum fitting algorithm, which is designed to robustly estimate the stellar continuum over the full wavelength range ($3850\,\text{\AA }\leq\lambda\leq 9000\,\text{\AA }$), and remains effective under varying SNRs. The method takes as input the observed wavelength array $\mathbf{w}$, the corresponding flux array $\mathbf{f}$, and an estimate of the average SNR, and returns a smooth continuum model $\mathbf{c}$.
@@
    Red Segment Correction: A fourth-order polynomial fit is applied iteratively to the red-side segment. At each step, outliers deviating more than $3\sigma$ below the fit (or any point below the fit) are replaced by the polynomial value. This process is repeated up to 8 iterations to suppress absorption features and stabilize the continuum estimate.
@@ Appendix C Normalizing Flows for Parameter Estimation
Normalizing flows provides a flexible and tractable approach to modeling complex probability distributions by transforming a simple base distribution through a sequence of invertible and differentiable mappings. In SBI, we use normalizing flows to learn an approximate posterior $q_{\phi}(\bm{\theta}\mid\mathbf{x})$, conditioned on observations $\mathbf{x}$, following the Neural Posterior Estimation (NPE) framework.
@@
$q_{\phi}(\bm{\theta}\mid\mathbf{x})=p_{Z}(f^{-1}_{\phi}(\bm{\theta};\mathbf{x}))\left|\det\left(\frac{\partial f^{-1}_{\phi}(\bm{\theta};\mathbf{x})}{\partial\bm{\theta}}\right)\right|=p_{Z}(\mathbf{z})\left|\det\left(\frac{\partial f_{\phi}(\mathbf{z};\mathbf{x})}{\partial\mathbf{z}}\right)\right|^{-1},$ (C1)
@@ C.1 Masked Autoregressive Flow (MAF)
MAF (2017arXiv170507057P)) models the forward transformation $\mathbf{z}\mapsto\bm{\theta}$ as a sequence of autoregressive operations. Each component of $\bm{\theta}$ is computed as:
$\theta_{i}=\mu_{i}(\mathbf{z}_{<i};\mathbf{x})+\sigma_{i}(\mathbf{z}_{<i};\mathbf{x})\cdot z_{i},$ (C4)
where $\mu_{i}$ and $\sigma_{i}$ are outputs of neural networks that depend on the previous components $\mathbf{z}_{<i}$ (or equivalently, the previous output components $\bm{\theta}_{<i}$) and the conditioning variable $\mathbf{x}$. The Jacobian of this transformation is lower triangular, which allows the log-determinant to be computed efficiently:
$\log\left|\det\left(\frac{\partial\bm{\theta}}{\partial\mathbf{z}}\right)\right|=\sum_{i}\log\sigma_{i}(\mathbf{z}_{<i};\mathbf{x}).$ (C5)
@@ C.2 Neural Spline Flow (NSF)
NSF (2019arXiv190604032D) generalizes MAF by replacing affine transformations with monotonic rational-quadratic splines. Each component is transformed as:
$\theta_{i}=\text{spline}_{i}(z_{i};\psi_{i}(\mathbf{z}_{<i},\mathbf{x})),$ (C6)
where $\psi_{i}$ are spline parameters (bin widths, heights, and derivatives) predicted by a neural network conditioned on $\mathbf{z}_{<i}$ and $\mathbf{x}$. The log-determinant of the Jacobian is given by:
@@
where each derivative term is efficiently computed from the analytical form of the spline. Note that while this autoregressive form is analogous to MAF, the default implementation based on the provided code is often the coupling layer variant of NSF (NSF-C), which is typically more efficient for density evaluation. Details on the coupling layer structure can be referred to in 2019arXiv190604032D.
@@ Appendix D Pre-trained Models
In this paper, we tried two different kinds of pre-trained models, one is the transformer-based model and the other is a MLP-based auto-encoder. Both kinds of networks have the same number of trainable parameters (42.7 million). We use a batch size of 64 per GPU. The models are optimized using AdamW (2017arXiv171105101L) with a learning rate of $1\times 10^{-5}$ and a weight decay of 0.1. The learning rate follows a cosine annealing schedule with linear warm-up.
@@ D.1 Transformer-Based Spectral Pre-trained Models
We describe two transformer models designed for the masked reconstruction of stellar spectra, based on 2024MNRAS.531.4990P, each tailored to the characteristics of a different input data set: LAMOST and Gaia XP.
@@ D.1.1 Masked Spectral Modeling for LAMOST Spectra
The input is linearly embedded and added to the learned positional embeddings. Then it passes through $N=6$ layers of standard transformer blocks with $H=6$ attention heads. The output is normalized and decoded through a linear projection.
@@
To train the model in a self-supervised fashion, we applied a chunk-based masking strategy. Given an input sequence of length $T^{\prime}$, we conceptually divide it into $M=6$ segments of equal-length and randomly select a contiguous piece of width $w=10$ within each segment to mask. For each chunk, its starting index is sampled uniformly from the allowable range within the segment. This ensures that the masked regions are distributed across the sequence and sufficiently separated.
Formally, let the $i$-th segment span the indices $s_{i},s_{i}+\ell]$, where $\ell=\lfloor T^{\prime}/M\rfloor$. Then the masked region for segment $i$ is:
@@
where $m_{t}=1$ if $t$ is masked and 0 otherwise.
LAMOST spectra benefit from local patterns and detailed features. Thus, using dense chunk-based masking and slicing helps the transformer leverage locality while preserving the global context. We train this model with 8 GPUs over a total of 128 epochs, a process that takes roughly 20 hours.
@@ D.1.2 Masked Spectral Modeling for Gaia XP Spectra
As in the LAMOST model, we standardize the spectrum and pre-pend two (mean and standard deviation) tokens:
$x^{\prime}_{0}=\mu,\quad x^{\prime}_{1}=\sigma$ (D5)
@@
We train this model with 8 GPUs and a total of 191 epochs, which is roughly 40 hours.
@@ D.2 MLP-based Autoencoder
For comparison with the transformer-based models, we also construct an MLP-based autoencoder to pre-train the stellar spectra. For both LAMOST LRS and Gaia XP, the network architectures are similar: an initial projection layer of shape $\text{input\_dim},\text{hidden\_dim}]$; the encoder consists of two MLP blocks, each with structure $\text{hidden\_dim},3\times\text{hidden\_dim},\text{hidden\_dim}]$; followed by a bottleneck layer of shape $\text{hidden\_dim},768]$. The decoder mirrors this structure to reconstruct the input spectra. For LAMOST LRS, we use $\text{input\_dim}=1462+1$ and $\text{hidden\_dim}=1245$; for Gaia XP, we use $\text{input\_dim}=343+2$ and $\text{hidden\_dim}=1290$. The former includes one additional (logarithmic) standard deviation as the first token, while the latter includes both the mean and the standard deviation as the first two tokens. We train the LRS model on 4 GPUs for 128 epochs, taking approximately 140 minutes, and the Gaia XP model 4 GPUs for 191 epochs, taking approximately 150 minutes.
@@ E.1 Projection Networks
For the projection networks, the LAMOST LRS model follows 2024MNRAS.531.4990P. We first obtain the output from a pre-trained LAMOST LRS model, then go through a cross-attention module with a learnable query vector. The cross-attention module has four attention heads with an output dimension of 768. Finally, we have an MLP layer with hidden features that have dimension $4\times 768$. We did not compress the information more in this projected network in order to investigate the information gain without compression. Therefore, the dimension of the final projected embedding is still 768. The number of trainable parameters is about 7.1 million.
@@
Each pathway includes a residual MLP block at the end with hidden dimension of $4\times 512$ and $4\times 256$, respectively.
@@ E.2 Decoders
For the cross-modal decoder of the CLIP-split model, we take only the shared representation as input. The decoder architecture depends on the output dimension, yielding approximately 12.5 and 3.9 million trainable parameters for the following two setups:
@@ Appendix F Loss Curves
This section presents the comparisons of the loss curves between the CLIP-pr and CLIP-split models, as shown in Figure 8. Although the CLIP-pr model achieves a lower CLIP loss during training, it exhibits a lower absolute cosine similarity in test pairs, as shown in Table 2. This discrepancy may arise from two factors. First, the geometry of the high-dimensional embedding spaces—768 dimensions in the CLIP-pr projected embedding versus 512 in the CLIP-split shared embedding—tends to yield lower cosine similarity in the former. Second, the CLIP loss focuses on relative alignment rather than absolute similarity. Additionally, the CLIP-pr model yields slightly lower cross-modal prediction losses. In contrast, the CLIP-split model converges more quickly in learning the reconstruction and achieves marginally lower reconstruction loss, likely due to its design of a non-shared projected embedding space.
Appendix G Summary of Key Hyper-parameters and Configurations
Table 6 summarizes the key hyper-parameters and configurations used across the main training stages. For complete details and implementation settings, please refer to Appendix D, Appendix E, and Section II.
Table 6: Summary of hype-parameters and configurations across main training stages.
Setting LRS MT XP OAE CLIP CLIP-r CLIP-p
Encoder / Layers 6 Self-Attn (6 heads) 2 MLP blocks enc.+dec. Cross-Attn (4 heads)+MLP Cross-Attn (4 heads)+MLP Cross-Attn (4 heads)+MLP
Latent dim. 768 768 768 768 768
[…]
Masking / Rate Chunk-based ($\sim$45%)
Optimizer AdamW AdamW AdamW AdamW AdamW
Learning rate $\times 10^{-5} 1×1$$\times 10^{-5}$ $\times 10^{-4} 1×1$$\times 10^{-4}$ $1\times 10^{-4}$
Weight decay 0.10 0.10 0.05 0.05 0.05
LR schedule Cosine + warm-up Cosine + warm-up Cosine + warm-up Cosine + warm-up Cosine + warm-up
Batch / GPU 64 64 1024 1024 1024
GPUs 8 4 8 8 8
Epochs / Time 128 / 20 h 191 / 2.5 h 110 /3 h 110 / 3 h 110 / 3 h
Loss fn. Masked MSE MSE $\mathcal{L}_{\text{CLIP}}$ $\mathcal{L}_{\text{CLIP}}+w_{\text{recon}}\mathcal{L}_{\text{recon}}$ $\mathcal{L}_{\text{CLIP}}+w_{\text{pred}}\mathcal{L}_{\text{pred}}$
Loss weights $w_{\text{recon}}=1.0$ $w_{\text{pred}}=1.0$
CLIP temp $\tau$ 15.5 15.5 15.5
Setting CLIP-pr CLIP-split Downstream MLP Downstream SBI (NPE) Encoder / Layers Cross-Attn (4 heads)+MLP 2 Cross-Attn branches 4 MLP [in,1024,512,64,1] 2 transforms, 60 hidden units each Latent dim. 768 512+256 Tokenization Same as CLIP Same as CLIP Raw spectra / embeddings Raw spectra / embeddings Masking / Rate Optimizer AdamW AdamW AdamW Adam Learning rate $\times 10^{-4} 1×1$$\times 10^{-4}$ $1\times 10^{-5}$ $5\times 10^{-4}$ Weight decay 0.05 0.05 $1\times 10^{-4}$ LR schedule Cosine + warm-up Cosine + warm-up ReduceLROnPlateau Batch / GPU 1024 1024 32 50 GPUs 8 8 1 1 Epochs / Time 110 / 3 h 110 / 3 h Max. 100 / <<10 min – / <<10 min Loss fn. $\mathcal{L}_{\text{CLIP}}+w_{\text{recon}}\mathcal{L}_{\text{recon}}+w_{\text{pred}}\mathcal{L}_{\text{pred}}$ $\mathcal{L}_{\text{CLIP}}+w_{\text{recon}}\mathcal{L}_{\text{recon}}+w_{\text{pred}}\mathcal{L}_{\text{pred}}$ MSE Negative log-likelihood Loss weights $w_{\text{recon}}=w_{\text{pred}}=1.0$ $w_{\text{recon}}=w_{\text{pred}}=1.0$ CLIP temp $\tau$ 15.5 15.5